Circular No. 8142

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003fa NEAR MCG +07-36-33 T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report that a spectrum (range 370–750 nm) of the variable star near MCG +07-36-33 (cf. IAUC 8140), obtained by P. Berlind on June 5.42 UT with the Mt. Hopkins 1.5-m telescope (+ FAST), shows it to be a type-Ia supernova well before maximum. The NASA/IPAC Extragalactic Database (NED) recession velocity for MCG +07-36-33 is 1800 km/s, which is inconsistent with the observed spectrum. Comparison with early-time spectra of SN 2001V indicates that the recession velocity of the supernova is 10000–15000 km/s and that it is 10–14 days before maximum.

SUPERNOVAE 2003em, 2003ev, 2003ey, AND 2003ez M. Hamuy and M. Phillips, Carnegie Observatories, report that a spec-

trum (range 370–900 nm) of SN 2003em (cf. IAUC 8137), obtained on June 4.42 UT with the Las Campanas Observatory Clay 6.5-m telescope (+ LDSS-2), shows it to be a type-Ia supernova, ∼ 40 days past maximum. A spectrum of SN 2003ev (cf. IAUC 8140), obtained on June 4.19, shows a relatively featureless red continuum with broad undulations. The main features are broad emissions near Hα and 457 nm, with a trough centered around 570 nm that is possibly due to He I and Na I D. These features are consistent with a highly reddened young type-Ic supernova, but a definitive classification will require additional observations. Superimposed narrow emission Hα emission from the host galaxy yields a redshift of 0.024. A spectrum of SN 2003ey (cf. IAUC 8141), obtained on June 4.27, shows it to be a type-Ia supernova, with strong resemblance to the spectrum of SN 1994D taken 5 days past maximum (the comparison yielding a redshift of ≈ 0.065 for SN 2003ey). A spectrum of SN 2003ez (cf. IAUC 8141), taken on June 4.13, shows it to bea type-Ia supernova, bearing strong resemblance to the spectrum of SN 1994D taken 3 days before maximum, although the Si II 635.5-nm feature in SN 2003ez is weaker; the minimum of this line yields an expansion velocity of −10200 km/s (by adopting the NED host-galaxy recession velocity of 14343 km/s).

H. Yamaoka, Kyushu University, reports that K. Ayani found SN 2003ez at R ∼ 17.4 on an image obtained with the Bisei Astronomical Observatory 1.01-m telescope on June 4.53 UT. Despite some saturation of the image, Yamaoka adds that SN 2003ez does not appear present on a Digitized Sky Survey image.

2003 June 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green